# Magnitude homology of real hyperplane arrangements

### 04 Apr 2026 v1

@@ 4.2. Geodesic magnitude homology and diagonal tope graphs
Let us determine the geodesic magnitude homology of arrangements.
@@
Let $\mathcal{A}$ be an arrangement. Then its geodesic magnitude homology is as follow
$MH_{k,\ell}^{geod}(\mathcal{A})\cong\bigoplus_{F\in\mathcal{F}(\mathcal{A}):\mathrm{rank}\mathcal{A}_{F}=k,\#\mathcal{A}_{F}=\ell}\mathbb{Z}^{\#\mathrm{Ch}(\mathcal{A}_{F})}$
@@
By Corollary 2.22,
$MH_{k,\ell}^{geod}(\mathcal{A})\cong\bigoplus_{a,b\in\mathrm{Ch}(\mathcal{A}):d(a,b)=\ell}\widetilde{H}_{k-2}(C_{*}(a,b)_{\mathcal{T}(\mathcal{A})}).$
By Proposition 2.34, the interval $(a,b)_{\mathcal{T}(\mathcal{A})}$ is homotopy equivalent to $S^{\mathrm{rank}\mathcal{A}_{F}-2}$ if $a,b]_{\mathcal{T}(\mathcal{A})}=\mathrm{Ch}(\mathcal{A})_{F}$ for some $F\in\mathcal{F}(\mathcal{A})$ and contractible otherwise. Therefore each $F\in\mathcal{F}(\mathcal{A})$ such that $k=\mathrm{rank}\mathcal{A}_{F}$ and $\ell=\#\mathcal{A}_{F}$ contribute to $MH_{k,\ell}^{geod}(\mathcal{A})$ a copy of $\mathbb{Z}^{\#\mathrm{Ch}(\mathcal{A})_{F}}$ and there is no other summand. ∎
Recall that $c^{X}=\#\mathrm{Ch}(\mathcal{A}^{X})$ and let $c_{X}=\#\mathrm{Ch}(\mathcal{A}_{X})$. We can give a combinatorial description of geodesic magnitude Betti number.
@@
The geodesic magnitude Betti number
$\beta_{k,\ell}^{geod}(\mathcal{A})=\sum_{X\in L_{k}(\mathcal{A}):\#\mathcal{A}_{X}=\ell}c^{X}c_{X}.$
In particular, geodesic magnitude homology is determined by $L(\mathcal{A})$.
@@
We know that the Boolean arrangement $Bl_{d}$ has tope graph $\mathcal{T}(Bl_{d})=Q_{d}$, which is diagonal (see Example 2.14). Suppose $\mathcal{A}$ is not Boolean, then $n=\#\mathcal{A}>r=\mathrm{rank}\mathcal{A}$. Therefore
$\beta_{r,n}(\mathcal{A})\geq\beta_{r,n}^{geod}(\mathcal{A})=\#\mathrm{Ch}(\mathcal{A}),$
where the equation is from Corollary 4.3. This means $\mathcal{A}$ is not diagonal. ∎
@@ 4.3. Magnitude homology with support
Consider a proper chain $\vec{x}=(x_{0},\ldots,x_{k})\in P_{k,\ell}(\mathcal{A})$, let us define $S(\vec{x})=S(x_{0},x_{1})\cup\cdots\cup S(x_{k-1},x_{k})$ as the set of hyperplanes crossed by $\vec{x}$. Notice that $S(\vec{x})\leq\ell(\vec{x})$. It is not difficult to verify that $x_{i}$ is smooth in $\vec{x}$, i.e. $d(x_{i-1},x_{i+1})=d(x_{i-1},x_{i})+d(x_{i},x_{i+1})$, is equivalent to $S(x_{i-1},x_{i+1})=S(x_{i-1},x_{i})\sqcup S(x_{i},x_{i+1})$. So if a proper chain $\vec{y}$ appears in the boundary of $\vec{x}$, then $S(\vec{x})=S(\vec{y})$. In other words, the boundary operator $\partial:MC_{k,\ell}(\mathcal{A})\to MC_{k-1,\ell}(\mathcal{A})$ preserves crossed hyperplanes. This motivated us to define the following.
@@
For arrangement $\mathcal{A}$ and a subarrangement $\mathcal{B}\subseteq\mathcal{A}$, we define the $\mathcal{B}$-supported magnitude chain group $MC_{k,\ell}(\mathcal{A};\mathcal{B})$ as the subgroup of $MC_{k,\ell}(\mathcal{A})$ generated by $P_{k,\ell}(\mathcal{A};\mathcal{B})=\{\vec{x}\in P_{k,\ell}(\mathcal{A})\mid S(\vec{x})=\mathcal{B}\}$. The above discussion shows that $MC_{*,\ell}(\mathcal{A};\mathcal{B})$ forms a subcomplex of $MC_{*,\ell}(\mathcal{A})$, called the $\mathcal{B}$-supported magnitude complex of $\mathcal{A}$. Its homology $MH_{*,\ell}(\mathcal{A};\mathcal{B})$ is called the $\mathcal{B}$-supported magnitude homology of $\mathcal{A}$.
Definition 4.6.
When $\mathcal{B}=\mathcal{A}$, we call $MC_{*,\ell}^{\circ}(\mathcal{A}):=MC_{*,\ell}(\mathcal{A};\mathcal{A})$ the length $\ell$ interior magnitude complex of $\mathcal{A}$ and $MH_{*,\ell}^{\circ}(\mathcal{A}):=MH_{*,\ell}(\mathcal{A};\mathcal{A})$ the length $\ell$ interior magnitude homology of $\mathcal{A}$. Let us also define the interior magnitude Betti number $\beta_{k,\ell}^{\circ}(\mathcal{A}):=\mathrm{rank}MH_{k,\ell}^{\circ}(\mathcal{A})$ and the length $\ell$ interior magnitude Euler characteristic $\chi_{\ell}^{\circ}(\mathcal{A}):=\sum_{k}(-1)^{k}\beta_{k,\ell}^{\circ}(\mathcal{A})$.
@@
For arrangement $\mathcal{A}$, we have the decomposition of magnitude chain groups
@@
inducing a decomposition of homology groups
@@
Furthermore $MC_{k,\ell}(\mathcal{A};\mathcal{B})=0$ and $MH_{k,\ell}(\mathcal{A};\mathcal{B})=0$ if $\ell<\#\mathcal{B}$.
4.4. Face decomposition of magnitude homology
We would like to further decompose the magnitdue complex/homology with support. To this end, we introduce another notation.
@@ 4.4. Face decomposition of magnitude homology
For arrangement $\mathcal{A}$ and a subarrangement $\mathcal{B}\subseteq\mathcal{A}$. Consider an induced subgraph $G$ of $\mathcal{T}(\mathcal{A})$. We define the localized $\mathcal{B}$-supported magnitude chain group $MC_{k,\ell}(G;\mathcal{B})$ as the subgroup of $MC_{k,\ell}(\mathcal{A};\mathcal{B})$ generated by chains entirely contained in $G$,
@@
We must show that the boundary operator $\partial$ stabilizes the subgroup. By Proposition 4.7, $\partial$ strictly preserves the set $\mathcal{B}$ of crossed hyperplanes. Furthermore, by Lemma 2.26, the $F$-subgraph $\mathcal{T}(\mathcal{A})_{F}$ is gated, and therefore convex (Proposition 2.2). Because the boundary operator $\partial_{i}$ merely deletes the chamber $x_{i}$ from the sequence, all remaining chambers in the chain $\partial_{i}\vec{x}$ are still contained within the vertex set of $\mathcal{T}(\mathcal{A})_{F}$. Thus, $\partial(MC_{k,\ell}(\mathcal{T}(\mathcal{A})_{F};\mathcal{B}))\subseteq MC_{k-1,\ell}(\mathcal{T}(\mathcal{A})_{F};\mathcal{B})$. ∎
Now we are ready to state the main structural result of the magnitude homology of an arrangement.
@@
For an arrangement $\mathcal{A}$, we have the face decomposition of its magnitude complex
@@
inducing the face decomposition of magnitude homology
@@
By Proposition 4.7, the global magnitude chain complex decomposes into a direct sum over all subarrangements $\mathcal{B}\subseteq\mathcal{A}$,
@@
For a fixed subarrangement $\mathcal{B}$, consider a chain $\vec{x}\in P_{k,\ell}(\mathcal{A};\mathcal{B})$. By Definition 4.5, $S(\vec{x})=\mathcal{B}$, meaning the chain never crosses any hyperplane in the complement $\mathcal{A}\setminus\mathcal{B}$. Therefore, every chamber in $\vec{x}$ shares a strictly constant, non-zero sign on all hyperplanes in $\mathcal{A}\setminus\mathcal{B}$. This constant sign pattern uniquely defines a face $F\in\mathcal{F}(\mathcal{A})$ such that $F_{H}=0$ for $H\in\mathcal{B}$, and $F_{H}\in\{+,-\}$ for $H\notin\mathcal{B}$. By definition, the subarrangement associated with this face is exactly $\mathcal{A}_{F}=\mathcal{B}$. Because the sign pattern on $\mathcal{A}\setminus\mathcal{B}$ is uniquely determined for each chain, the subgroup $MC_{k,\ell}(\mathcal{A};\mathcal{B})$ further decomposes into a direct sum over the faces $F$ that satisfy $\mathcal{A}_{F}=\mathcal{B}$,
$MC_{k,\ell}(\mathcal{A};\mathcal{B})=\bigoplus_{F\in\mathcal{F}(\mathcal{A}):\mathcal{A}_{F}=\mathcal{B}}MC_{k,\ell}(\mathcal{T}(\mathcal{A})_{F};\mathcal{B}).$
By Lemma 2.26, the $F$-subgraph $\mathcal{T}(\mathcal{A})_{F}$ is a gated subgraph isometrically isomorphic to the tope graph of the localized subarrangement $\mathcal{A}_{F}$. Under this isomorphism, the $\mathcal{B}$-supported complex maps identically to the full interior complex of $\mathcal{A}_{F}$. Therefore, by Definition 4.6,
$MC_{k,\ell}(\mathcal{T}(\mathcal{A})_{F};\mathcal{B})\cong MC_{k,\ell}(\mathcal{A}_{F};\mathcal{A}_{F})=MC^{\circ}_{k,\ell}(\mathcal{A}_{F}).$
Substituting this isomorphism back into the original sum over subarrangements $\mathcal{B}$ yields the result
$MC_{k,\ell}(\mathcal{A})\cong\bigoplus_{\mathcal{B}\subseteq\mathcal{A}}\left(\bigoplus_{F:\mathcal{A}_{F}=\mathcal{B}}MC^{\circ}_{k,\ell}(\mathcal{A}_{F})\right)=\bigoplus_{F\in\mathcal{F}(\mathcal{A})}MC^{\circ}_{k,\ell}(\mathcal{A}_{F}).$
Since this is a decomposition into subcomplexes, we have the face decomposition of magnitude homology
For an arrangement $\mathcal{A}$, let us write
$\chi^{\circ}(\mathcal{A},q):=\sum_{\ell=0}^{\infty}\chi_{\ell}^{\circ}(\mathcal{A})q^{\ell}\in\mathbb{Z}[\![q]\!].$
Then we have a reformulation of Conjecture 3.7.
@@
Let $r=\mathrm{rank}\mathcal{A}$ and $n=\#\mathcal{A}$. The following are equivalent
  1. (1)
    $\mathrm{Mag}(\mathcal{A},q)=\sum_{F}\mathrm{Mag}^{\circ}(\mathcal{A}_{F},q)$;
  2. (2)
    $\chi^{\circ}(\mathcal{A},q)=\mathrm{Mag}^{\circ}(\mathcal{A},q)$;
  3. (3)
    $\chi_{\ell+n}^{\circ}(\mathcal{A})=(-1)^{r}\chi_{\ell}(\mathcal{A})$ for $\ell\geq 0$.
@@
Since Euler characteristic is additive over direct sums, the face decomposition of magnitude homology (Theorem 4.10) yields
$\chi_{\ell}(\mathcal{A})=\sum_{F\in\mathcal{F}(\mathcal{A})}\chi_{\ell}^{\circ}(\mathcal{A}_{F}).$
Multiplying by $q^{\ell}$ and summing over $\ell$, we obtain
$\mathrm{Mag}(\mathcal{A},q)=\sum_{F\in\mathcal{F}(\mathcal{A})}\chi^{\circ}(\mathcal{A}_{F},q),$
By Möbius inversion on $\mathcal{F}(\mathcal{A})$, we see that (1) is equivalent to (2). The equivalence between (2) and (3) is straightforward by definition. ∎
@@
Note that if there is a (quasi-)isomorphism of chain complexes
$MC_{*,\ell}(\mathcal{A})\to MC_{*+r,\ell+n}^{\circ}(\mathcal{A})=MC_{*+r,\ell+n}(\mathcal{A};\mathcal{A})$
for all $\ell\geq 0$, then the conditions in Corollary 4.11 hold.

### 01 May 2026 v3

@@ 5.2. Geodesic magnitude homology and diagonal tope graphs
We now determine the geodesic magnitude homology of arrangements. Recall that $c^{X}=\#\mathrm{Ch}(\mathcal{A}^{X})$. Let $c_{X}=\#\mathrm{Ch}(\mathcal{A}_{X})$. Both $c^{X}$ and $c_{X}$ are determined by the intersection lattice $L(\mathcal{A})$.
@@
Let $\mathcal{A}$ be an arrangement. Then its geodesic magnitude homology is given as follows:
$MH_{k,\ell}^{\mathrm{geod}}(\mathcal{A})\cong\bigoplus_{\begin{subarray}{c}F\in\mathcal{F}(\mathcal{A})\\ \operatorname{rank}\mathcal{A}_{F}=k,\ \#\mathcal{A}_{F}=\ell\end{subarray}}\mathbb{Z}^{\#\mathrm{Ch}(\mathcal{A}_{F})}\cong\bigoplus_{\begin{subarray}{c}X\in L_{k}(\mathcal{A})\\ \#\mathcal{A}_{X}=\ell\end{subarray}}\mathbb{Z}^{c^{X}c_{X}}.$
In particular, $MH_{k,\ell}^{\mathrm{geod}}(\mathcal{A})$ is determined by $L(\mathcal{A})$.
@@
By Corollary 2.24, we have
$MH_{k,\ell}^{\mathrm{geod}}(\mathcal{A})\cong\bigoplus_{\begin{subarray}{c}a,b\in\mathrm{Ch}(\mathcal{A})\\ d(a,b)=\ell\end{subarray}}\widetilde{H}_{k-2}(C_{*}(a,b)_{\mathcal{T}(\mathcal{A})}).$
By Proposition 2.35, the interval $(a,b)_{\mathcal{T}(\mathcal{A})}$ is homotopy equivalent to $S^{\operatorname{rank}\mathcal{A}_{F}-2}$ if $a,b]_{\mathcal{T}(\mathcal{A})}=\mathrm{Ch}(\mathcal{A})_{F}$ for some $F\in\mathcal{F}(\mathcal{A})$, and is contractible otherwise. Therefore each $F\in\mathcal{F}(\mathcal{A})$ such that $k=\operatorname{rank}\mathcal{A}_{F}$ and $\ell=\#\mathcal{A}_{F}$ contributes a copy of $\mathbb{Z}^{\#\mathrm{Ch}(\mathcal{A})_{F}}$ to $MH_{k,\ell}^{\mathrm{geod}}(\mathcal{A})$, and there is no other summand. ∎
@@
Let $\mathcal{A}$ be an arrangement. Then its geodesic magnitude Betti numbers are given by
$\beta_{k,\ell}^{\mathrm{geod}}(\mathcal{A})=\sum_{\begin{subarray}{c}X\in L_{k}(\mathcal{A})\\ \#\mathcal{A}_{X}=\ell\end{subarray}}c^{X}c_{X}.$
@@
We know that the Boolean arrangement $Bl_{d}$ has tope graph $\mathcal{T}(Bl_{d})=Q_{d}$, which is diagonal (see Example 2.14). Suppose $\mathcal{A}$ is not Boolean. Then $n=\#\mathcal{A}>r=\operatorname{rank}\mathcal{A}$. We have
$\beta_{r,n}(\mathcal{A})\geq\beta_{r,n}^{\mathrm{geod}}(\mathcal{A})=\#\mathrm{Ch}(\mathcal{A}),$
where the equality follows from Corollary 5.3. This means that $\mathcal{A}$ is not diagonal. ∎
@@ 5.3. Magnitude homology with support
Consider a proper chain $\vec{x}=(x_{0},\ldots,x_{k})\in P_{k,\ell}(\mathcal{A})$, and define $S(\vec{x})=S(x_{0},x_{1})\cup\cdots\cup S(x_{k-1},x_{k})$ as the set of hyperplanes crossed by $\vec{x}$. Observe that $S(\vec{x})\leq\ell(\vec{x})$. One verifies that $x_{i}$ is smooth in $\vec{x}$, i.e. $d(x_{i-1},x_{i+1})=d(x_{i-1},x_{i})+d(x_{i},x_{i+1})$, if and only if $S(x_{i-1},x_{i+1})=S(x_{i-1},x_{i})\sqcup S(x_{i},x_{i+1})$. Thus, if a proper chain $\vec{y}$ appears in the boundary of $\vec{x}$, then $S(\vec{x})=S(\vec{y})$. In other words, the boundary operator $\partial:MC_{k,\ell}(\mathcal{A})\to MC_{k-1,\ell}(\mathcal{A})$ preserves crossed hyperplanes. This motivates the following definition.
@@
For an arrangement $\mathcal{A}$ and a subarrangement $\mathcal{B}\subseteq\mathcal{A}$, we define the $\mathcal{B}$-supported magnitude chain group $MC_{k,\ell}(\mathcal{A};\mathcal{B})$ as the subgroup of $MC_{k,\ell}(\mathcal{A})$ generated by $P_{k,\ell}(\mathcal{A};\mathcal{B})=\{\vec{x}\in P_{k,\ell}(\mathcal{A})\mid S(\vec{x})=\mathcal{B}\}$. The above discussion shows that $MC_{*,\ell}(\mathcal{A};\mathcal{B})$ forms a subcomplex of $MC_{*,\ell}(\mathcal{A})$, called the $\mathcal{B}$-supported magnitude complex of $\mathcal{A}$. Its homology $MH_{*,\ell}(\mathcal{A};\mathcal{B})$ is called the $\mathcal{B}$-supported magnitude homology of $\mathcal{A}$.
@@
For an arrangement $\mathcal{A}$, we have the following decomposition of magnitude chain groups:
@@
which induces a decomposition of homology groups
@@
Furthermore, $MC_{k,\ell}(\mathcal{A};\mathcal{B})=MH_{k,\ell}(\mathcal{A};\mathcal{B})=0$ if $\ell<\#\mathcal{B}$.
Using the notion of magnitude homology with support, we define interior magnitude homology. Later, we show that this is a categorification of the interior magnitude defined in Definition 3.6.
@@
Let $\mathcal{A}$ be an arrangement and let $X\in L(\mathcal{A})$ be a flat. We define the $X$-supported magnitude complex of $\mathcal{A}$ by
$MC_{*,\ell}(\mathcal{A};X):=\bigoplus_{\begin{subarray}{c}\mathcal{B}\subseteq\mathcal{A}\\ \cap\mathcal{B}=X\end{subarray}}MC_{*,\ell}(\mathcal{A};\mathcal{B}).$
Its homology $MH_{*,\ell}(\mathcal{A};X)$ is called the $X$-supported magnitude homology of $\mathcal{A}$. In particular, we define
$MC_{*,\ell}^{\circ}(\mathcal{A}):=MC_{*,\ell}(\mathcal{A};\cap\mathcal{A}),$
and call it the length $\ell$ interior magnitude complex of $\mathcal{A}$. We call $MH_{*,\ell}^{\circ}(\mathcal{A}):=H_{*}(MC_{*,\ell}^{\circ}(\mathcal{A}))$ the length $\ell$ interior magnitude homology of $\mathcal{A}$. We also define the interior magnitude Betti number $\beta_{k,\ell}^{\circ}(\mathcal{A}):=\operatorname{rank}MH_{k,\ell}^{\circ}(\mathcal{A})$ and the length $\ell$ interior magnitude Euler characteristic $\chi_{\ell}^{\circ}(\mathcal{A}):=\sum_{k}(-1)^{k}\beta_{k,\ell}^{\circ}(\mathcal{A})$.
5.4. Face decomposition for magnitude homology
Here we prove a homological version of the face decomposition formula established in Theorem 3.7. To this end, we introduce another notation.
Definition 5.8.
Let $\mathcal{A}$ be an arrangement and let $\mathcal{B}\subseteq\mathcal{A}$ be a subarrangement. Consider an induced subgraph $G$ of $\mathcal{T}(\mathcal{A})$. We define the localized $\mathcal{B}$-supported magnitude chain group $MC_{k,\ell}(G;\mathcal{B})$ as the subgroup of $MC_{k,\ell}(\mathcal{A};\mathcal{B})$ generated by chains entirely contained in $G$:
@@ 5.4. Face decomposition for magnitude homology
We must show that the boundary operator $\partial$ stabilizes the subgroup. By Proposition 5.6, $\partial$ strictly preserves the set $\mathcal{B}$ of crossed hyperplanes. Furthermore, by Lemma 2.29, the $F$-subgraph $\mathcal{T}(\mathcal{A})_{F}$ is gated, and therefore convex (Proposition 2.2). Because the boundary operator $\partial_{i}$ only deletes the chamber $x_{i}$ from the sequence, all remaining chambers in the chain $\partial_{i}\vec{x}$ are still contained in the vertex set of $\mathcal{T}(\mathcal{A})_{F}$. Thus, $\partial(MC_{k,\ell}(\mathcal{T}(\mathcal{A})_{F};\mathcal{B}))\subseteq MC_{k-1,\ell}(\mathcal{T}(\mathcal{A})_{F};\mathcal{B})$. ∎
Proposition 5.10.
Let $\mathcal{A}$ be an arrangement and let $X\in L(\mathcal{A})$ be a flat. Then the $X$-supported magnitude complex can be decomposed as follows:
$MC_{*,\ell}(\mathcal{A};X)\cong\bigoplus_{\begin{subarray}{c}F\in\mathcal{F}(\mathcal{A})\\ s(F)=X\end{subarray}}MC_{*,\ell}^{\circ}(\mathcal{A}_{F}).$
Proof.
By Definition 5.7, $MC_{*,\ell}(\mathcal{A};X)=\bigoplus_{\mathcal{B}\subseteq\mathcal{A},\cap\mathcal{B}=X}MC_{*,\ell}(\mathcal{A};\mathcal{B})$. For a subarrangement $\mathcal{B}$ with $\cap\mathcal{B}=X$, consider a chain $\vec{x}\in P_{k,\ell}(\mathcal{A};\mathcal{B})$. By Definition 5.5, $S(\vec{x})=\mathcal{B}$. The chain $\vec{x}$ never crosses any hyperplane in $\mathcal{A}\setminus\mathcal{A}_{X}$. Therefore, every chamber in $\vec{x}$ has a constant nonzero sign on each hyperplane $H\in\mathcal{A}\setminus\mathcal{A}_{X}$. These signs determine a chamber $F\in\mathrm{Ch}(\mathcal{A}^{X})$, or equivalently a face $F\in\mathcal{F}(\mathcal{A})$ with $s(F)=X$. Therefore the chain $\vec{x}$ lies in the $F$-subgraph $\mathcal{T}(\mathcal{A})_{F}$. By Lemma 2.29, $\mathcal{T}(\mathcal{A})_{F}$ is a gated subgraph of $\mathcal{T}(\mathcal{A})$ isometrically isomorphic to $\mathcal{T}(\mathcal{A}_{F})$. This induces a direct-sum decomposition of the $\mathcal{B}$-supported chains over the faces supported on $X$:
$MC_{k,\ell}(\mathcal{A};\mathcal{B})\cong\bigoplus_{\begin{subarray}{c}F\in\mathcal{F}(\mathcal{A})\\ s(F)=X\end{subarray}}MC_{k,\ell}(\mathcal{A}_{F};\mathcal{B}).$
Summing both sides over $\mathcal{B}\subseteq\mathcal{A}$ such that $\cap\mathcal{B}=X$, we obtain
$MC_{*,\ell}(\mathcal{A};X)\cong\bigoplus_{\begin{subarray}{c}F\in\mathcal{F}(\mathcal{A})\\ s(F)=X\end{subarray}}\bigoplus_{\begin{subarray}{c}\mathcal{B}\subseteq\mathcal{A}_{F}\\ \cap\mathcal{B}=X\end{subarray}}MC_{*,\ell}(\mathcal{A}_{F};\mathcal{B}).$
Since we are summing over $F$ supported on $X$, we have $\mathcal{A}_{F}=\mathcal{A}_{X}$ and $\cap\mathcal{A}_{F}=X$, so the inner sum is exactly $MC_{*,\ell}^{\circ}(\mathcal{A}_{F})$. This proves the desired formula. ∎
We are now ready to state the main structural result for the magnitude homology of an arrangement.
@@
For an arrangement $\mathcal{A}$, we have the following face decomposition of its magnitude complex:
@@
which induces the following face decomposition of magnitude homology:
@@
By Proposition 5.6, the magnitude chain complex decomposes into a direct sum over all subarrangements $\mathcal{B}\subseteq\mathcal{A}$,
@@
Grouping the $\mathcal{B}$-summands according to $\cap\mathcal{B}$, we have
$MC_{k,\ell}(\mathcal{A})=\bigoplus_{X\in L(\mathcal{A})}\bigoplus_{\begin{subarray}{c}\mathcal{B}\subseteq\mathcal{A}\\ \cap\mathcal{B}=X\end{subarray}}MC_{k,\ell}(\mathcal{A};\mathcal{B})=\bigoplus_{X\in L(\mathcal{A})}MC_{k,\ell}(\mathcal{A};X).$
By Proposition 5.10, the $X$-supported complex decomposes into interior complexes
$MC_{k,\ell}(\mathcal{A})\cong\bigoplus_{X\in L(\mathcal{A})}\bigoplus_{\begin{subarray}{c}F\in\mathcal{F}(\mathcal{A})\\ s(F)=X\end{subarray}}MC_{k,\ell}^{\circ}(\mathcal{A}_{F})=\bigoplus_{F\in\mathcal{F}(\mathcal{A})}MC_{k,\ell}^{\circ}(\mathcal{A}_{F}).$
The decomposition of homology follows immediately. ∎
@@
Let $\mathcal{A}$ be an arrangement with $\operatorname{rank}\mathcal{A}=r$ and $\mathcal{A}=n$. Then we have
$\mathrm{Mag}^{\circ}(\mathcal{A})=\sum_{\ell=0}^{\infty}\chi_{\ell}^{\circ}(\mathcal{A})q^{\ell}.$
Moreover, we have $\chi_{\ell+n}^{\circ}(\mathcal{A})=(-1)^{r}\chi_{\ell}(\mathcal{A})$ for $\ell\geq 0$, and $\chi_{\ell}^{\circ}(\mathcal{A})=0$ for $\ell<n$.
@@
Define $\Delta(\mathcal{A})$ to be the difference of the two expressions:
$\Delta(\mathcal{A})=\mathrm{Mag}^{\circ}(\mathcal{A})-\sum_{\ell=0}^{\infty}\chi_{\ell}^{\circ}(\mathcal{A})q^{\ell}.$
We prove $\Delta(\mathcal{A})=0$ by induction on the rank of $\mathcal{A}$. By Theorem 3.7 and Theorem 5.11, we have
$\mathrm{Mag}(\mathcal{A})=\sum_{F\in\mathcal{F}(\mathcal{A})}\mathrm{Mag}^{\circ}(\mathcal{A}_{F}),\quad\chi_{\ell}(\mathcal{A})=\sum_{F\in\mathcal{F}(\mathcal{A})}\chi^{\circ}_{\ell}(\mathcal{A}_{F}).$
Since $\mathrm{Mag}(\mathcal{A})$ can be written as $\sum_{\ell}\chi_{\ell}(\mathcal{A})q^{\ell}$, we obtain
$\sum_{F\in\mathcal{F}(\mathcal{A})}\Delta(\mathcal{A}_{F})=0.$
The induction hypothesis shows that $\Delta(\mathcal{A}_{F})=0$ for every face $F$ other than $\bigcap\mathcal{A}$. Thus $\Delta(\mathcal{A})=0$. The final assertion follows from the identity $\mathrm{Mag}^{\circ}(\mathcal{A})=(-1)^{r}q^{n}\mathrm{Mag}(\mathcal{A})$. ∎
@@
The identity $\chi_{\ell+n}^{\circ}(\mathcal{A})=(-1)^{r}\chi_{\ell}(\mathcal{A})$ in Corollary 5.12 is analogous to the Ehrhart–Macdonald reciprocity theorem for lattice polytopes [11, 26] and Stanley’s reciprocity theorem for posets [31].
Remark 5.14.
In light of Corollary 5.12, it is natural to expect that there exists an isomorphism of homology groups
$MH_{k,\ell}(\mathcal{A})\cong MH_{k+r,\ell+n}^{\circ}(\mathcal{A})$
for all $k,\ell\geq 0$, and that the interior magnitude homology $MH_{k,\ell}^{\circ}(\mathcal{A})$ vanishes for $0\leq\ell<n$.